100vt Aur 220 vlt ka electric current aur resistance kya hoga

Grade:12

1 Answers

Arun
5 years ago
Dear Ishan
 
P = Vi
100 = 220 * i
 
i = 100/220 = 0.455
 
and V = i R
hence R = V/i = 220/0.455 = 483.516 ohm